# What’s DAT? Three Case Studies of Measuring Software Development Productivity at Meta With Diff Authoring Time

### v1

@@ Abstract.
This paper introduces Diff Authoring Time (DAT), a powerful, yet conceptually simple approach to measuring software development productivity that enables rigorous experimentation. DAT is a time-based metric, which assess how long engineers take to develop changes, using a privacy-aware telemetry system integrated with version control, the IDE, and the OS. We validate DAT through observational studies, surveys, visualizations, and descriptive statistics. At Meta, DAT has powered experiments and case studies on more than 20 projects. Here, we highlight (1) an experiment on introducing mock types (a 14% DAT improvement), (2) the development of automatic memoization in the React compiler (33% improvement), and (3) an estimate of thousands of DAT hours saved annually through code sharing ($>50$% improvement). DAT offers a precise, yet high-coverage measure for development productivity, aiding business decisions. It enhances development efficiency by aligning the internal development workflow with the experiment-driven culture of external product development. On the research front, DAT has enabled us to perform rigorous experimentation on long-standing software engineering questions such as “do types make development more efficient?”
@@
booktitle: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering (FSE ’25), June 23–27, 2025, Trondheim, Norwaycopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM International Conference on the Foundations of Software Engineering (FSE); June 23rd-27th, 2025; Trondheim, NGisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineeringccs: Software and its engineering Software development methods
@@ 1. Introduction
For nearly two decades, Continuous Experimentation (CE) has been integral to external-facing product development in Big Tech (Kohavi et al., 2009; Tang et al., 2010; Karrer et al., 2021): Engineers frequently launch A/B experiments to evaluate new features, ensuring measurements are free of interferences. With a large user base, even small effect sizes are detectable. In contrast, internal development of developer tools used by thousands of developers mostly still relies on basic usage numbers and qualitative feedback (Forsgren et al., 2024). This is puzzling given the high value of engineering time, a major operational expense. To improve internal development, a metrics-driven culture is needed. However, measuring productivity accurately is challenging (Jaspan and Sadowski, 2019; Murphy-Hill et al., 2019; Beller et al., 2020), as traditional metrics are often inaccurate and easily gameable.
In this paper, we propose a solution inspired by Big Tech’s experiment-driven product culture. We introduce Diff Authoring Time (DAT) as a key metric for assessing development, not developer, productivity. DAT measures the time engineers take to develop a change, using a privacy-aware telemetry system integrated with version control, the IDE, and the operating system. We validated DAT through observational studies, surveys, visualizations, and statistics. Out of more than 20 studies completed with DAT, we highlight three, demonstrating DAT’s impact on productivity:
@@
    Introduction of mock types in Hack led to a 14% DAT improvement, marking a first in linking language-level features to productivity gains.
@@
    Quantification of productivity gains from a new feature in the React Compiler, auto-memoization, resulting in 33% productivity wins.
@@
    Development of a method to estimate efficiency gains from code sharing, showing a $>50$% improvement across frameworks, saving thousands of hours annually.
2. Background
Figure 1. Example illustration of the most relevant DAT-covered and uncovered activities for hypothetical diffs D123 and D987, and a comparison with other related metrics Code Gestation Time (CGT) and Time Spent by Diff (TSD).
@@ 2.1. Continuous Experimentation
The phrase “Continuous Experimentation” (CE) refers to the practice of regularly running experiments to validate one’s hypotheses, often simultaneously. While this approach has become widespread in large, software-intense organizations and has defined much of the engineering work in Big Tech (Kohavi et al., 2009), it is not ubiquitous (Mäntylä et al., 2022). Challenges include a lack of resources and knowledge on CE, particularly in startups.
@@
Melegati et al. establish a “CE maturity model” with five stages of adoption and implementation (Melegati et al., 2020). Before this paper, Infra was at a similar stage to a typical startup: while Meta as a whole was naturally aware of CE (stage 1), it required a cultural mind shift to embrace it for Infra (stage 2). An organic need driven by lower observed organizational output combined with a top-down mandate initiated this. To instill the new mindset and work processes throughout the organization, a particularly rigorous validation (cf. Section 4) was necessary (stage 3). We derived results by focusing on a generally accepted measure (DAT), and integrating it into our already validated experiment infrastructure for product (stage 4). Highly skilled data scientists performed much of this work, leading to fast productionization of these results, including publishing the achieved wins in this paper (stage 5+).
2.2. Big Tech and organizational structure
By “Big Tech”, we refer to major software companies such as Alphabet, Apple, Microsoft, and Meta, with billions of users (Fernandez et al., 2020). Their internal development teams often support tens of thousands of engineers (Hedlund, 2021), optimizing tools and workflows to ultimately boost company-wide productivity.
While experimentation only makes sense given a large enough developer base, we believe DAT can apply to a multitude of stakeholders outside of Big Tech. Power analysis can reveal that sometimes a few dozen developers could be sufficient to detect effect sizes of the magnitude we present in this paper (Fitzner and Heckinger, 2010). It is also not strictly necessary to have full control over the complete workflow, as one is typically interested in one specific aspect when running experiments. Many consultancies, tool vendors, open source projects, and even some researchers (Beller et al., 2017) attract equally large or larger developer user bases, with the Eclipse IDE for Java surpassing 12 million monthly downloads in 2023, for example (Krill, 2023). Moreover, even in the absence of so many users, can the mental model behind DAT help shape how we as developers engineer with a data-driven mindset.
2.3. Tools and Processes at Meta
A diff represents a single code change including a natural text title, summary, and test plan (Beller et al., 2023), similar to a pull request (Gousios et al., 2014). Almost all code-based changes end up as (ideally, self-contained) diffs. A typical way of creating code diffs is by creating commits in an IDE such as VS Code (Marcey, 2019), using Sapling, Meta’s integrated, platform-independent version control client for multiple version control systems (Goode and Agarwal, 2014; Durham Goode, 2022). However, there are a number of other modalities for creating diffs, including bespoke tools (Murali et al., 2024). Code review of all diffs happens on a custom version of the open-source Phabricator tool (Kudrjavets et al., 2022; Beller et al., 2023). Overall, these characteristics make diffs a natural and high-coverage measurement unit for productivity and experimentation purposes.
Many productivity experiments make sense on the diff level, meaning that one developer can appear both in the A and B group, naturally removing one of the biggest confounders. For example, a diff-level experiment could be to test for the DAT effects of a new Continuous Integration (CI) signal, which would not affect the developer’s core workflow or state of mind if absent in the future, as one among many hundreds of CI signals. Similarly, switching out the base language model from LLaMA-33B to LLaMA-65B (Touvron et al., 2023) would likely be an opaque operation to a developer. Other experiments, however, (additionally) need to be on the cohort level—it makes little sense and is practically impossible to switch one developer back-and forth between two VS Code versions with completely different user interfaces, for example. A diff-based measurement is still essential in these cases, as the diff level provides normalization for many other confounders such as length, complexity, language, framework, or testing.
2.4. Related Metrics
Productivity metrics in Software Engineering have a long history (Brooks, 1974), with various metrics proposed, such as lines of code (Jones and Nelson, 1976), change count (Meli, 2001), and perceived productivity (Meyer et al., 2017). However, these metrics are often flawed, gameable, subjective, impossible to obtain accurately, or some combination of the above (Kitchenham, 1997; Petersen, 2011; Symons, 1988; Huijgens and Solingen, 2014). Some even suggest abandoning productivity measurement altogether (Ko, 2019).
@@ 2.4. Related Metrics
Productivity frameworks such as SPACE (Forsgren et al., 2021) and DX Core4 (Noda et al., 2024) have recently provided frameworks on how to think about software development productivity at large, deferring the exact metric choice to their users. The metrics that are then often used tend to focus on easily measurable aspects of the development process, rather than the complexity of the authoring. For example, output metrics (such as lines of code, change count, …) emphasize the result of development while failing to capture the effort and complexity involved in the authoring lifecycle. Similarly, review time metrics, while good for measuring collaboration and delayed reviews, are not a measure of authoring time.
@@
Therefore, there was a need to develop and test a metric that captured the complexity of authoring: DAT. DAT does this by measuring the active time taken to develop a diff while enabling insight into the components that impact it, such as IDE features and programming languages. For instance, feature experimentation using DAT enables teams to assess the impact of specific improvements on developer velocity.
@@
We propose DAT as a conceptually simple, yet hard to implement, maximally objective metric that is not merely theoretical, but of high practical value. DAT supersedes Time Spent by Diff (TSD) and Code Gestation Time (CGT), illustrated in Figure 1. TSD averages coding time in a given period by the number of diffs published in that period, leading to misattribution and lacking diff-level precision. CGT measured time from coding start to diff landing, mixing active engineering with release time, complicating interpretation and lacking validation.
We adopt a pragmatic approach, focusing on accurate time measurement, as “time spent coding” aligns with developers’ productivity perceptions (Beller et al., 2020) and can be measured at scale.
@@ 3.1. High-level description and vision
Diff Authoring Time (DAT) aims to capture all active human developer time spent on a particular diff, split by the different roles involved (e.g., author and reviewer), as depicted in Figure 1. It aims to make this information available in a privacy-aware manner. The key characteristic of DAT is the accurate representation of time spent in the coding editor, an Integrated Development Environment (IDE). DAT builds upon this to cover all diff-related activities, tools, and roles. Naturally, time off the computer, such as sketching whiteboarding or meeting time will always be impossible or difficult to track.
@@
While there is an ongoing scientific debate on humans’ ability to effectively multitask (Koch et al., 2018; Skaugset et al., 2016), active human work on a computer tends to be highly sequential, even if such individual sequences can be extremely short (e.g., rapid application switching often happens in $<1s$): there can only be one mouse pointer, one cursor, and one application in focus at a given time, and such activity is normally related to one diff only.
@@
This gives DAT some advantageous properties that lend themselves to validation: it is non-overlapping (per user), meaning that two diffs by the same author must not have any overlapping time. As a corollary, DAT cannot exceed 24 hours per day (though it is, in practice, much shorter), can sensibly be aggregated (as opposed to CGT), and the sum of DAT in one day can be bounded by TSD. In fact, the comparison with TSD can reveal (1) improvement potential for DAT’s matching algorithm or (2) potential for efficiency improvement, as coding-related activities did not result in productive output.
The goal for DAT is to be used purely as a productivity metric for tooling and workflows, not people. As a result, DAT and its underlying telemetry were designed with privacy in mind.
3.2. Core DAT algorithm: Precise matches
Figure 2. Explanation of the core DAT commit matching algorithm.
@@ 3.2. Core DAT algorithm: Precise matches
As the first two rows in Figure 1 depict, the core authoring of a diff starts with the initial trace of an activity that we can connect with this diff. Typically, this is the start of the first IDE session—most often in VS Code—in which the first Sapling commit is made that is then later connected to that diff. DAT ends with the final IDE session connected to the diff. DAT includes only the time actively working on this diff in-between the two timestamps. DAT is effectively a non-overlapping function of a diff, meaning that DAT(D123) $\cap$ DAT(D987) = $\emptyset$. We further split DAT into an author and reviewer time component. The core DAT algorithm prioritises precise, high-confidence matching—it omits IDE activity that does not result in a commit, or activity for commits that do not result in a diff. It only operates on tools that allow such high confidence attribution: IDEs and the Phabricator Diff review tool.
We obtain basic activity and foreground application through the privacy-aware OS-level telemetry. We then correlate it to commits using telemetry from the Sapling plugin. Sapling has the advantage of tracking the underlying repository, meaning that any interaction with it (e.g., also from the command line) is captured.
@@
Figure 2 depicts the fundamental principle of DAT’s diff-to-session matching algorithm using the example of two diffs D123 and D987 that are developed sequentially in the same VS Code session: on top, the figure shows human activity as work sessions, along with point-in-time interactions with version control. It also shows the state of the repository (the output of sl status, similar to git status)—the current commit tracked by the version control system. The figure depicts the principle that any commit $CH_{x}$ is attributed on the time axis to the IDE session $s_{(x-1)}$ that precedes it. Thus, at its core, the DAT algorithm performs a 1-left-shift of the underlying commits with regard to the telemetry sessions. For example, the commit $CH1$ made at time $t2$ is attributed to the VS Code session $s1$ that happened from $t1$ to $t2.$ This is because any change developed starting from $t1$ would become part of the commit $CH1$. Similarly, both $s2$ and $s4$ are mapped to $CH2$. $s2$ is interrupted by the offline activity $s3$, which does not count toward DAT.
@@
There are two corner cases: (1) Any time a developer opens VS Code, an initial automatic checkout happens ($CH0$ in Figure 2). As the developer does not modify this pre-existing commit, the commit event from this automatic checkout is left-shifted away by the DAT algorithm. Similarly, DAT removes the later manual checkout of the pre-existing commit $CH7$ via explicit filtering. The commit $CH8$, which becomes $CH7$’s successor, is equally attributed to D987. (2) For the “trailing session” $s7$, there is no commit mapping because any work done there would not actually be saved as a commit.
@@
We now have a precise commit-to-session mapping (hence the algorithm name “precise matches”), but not yet a mapping to diffs. This, though, is trivial to achieve, as Version Control provides us with a list of all commits that make up a diff.
3.3. Additional DAT Heuristic: Anchor sessions

### v2

@@ Abstract.
This paper introduces Diff Authoring Time (DAT), a precise, yet conceptually simple approach to measuring software development productivity that enables rigorous experimentation. DAT is a time-based metric, which assesses how much active work time engineers take to develop self-contained changes. It uses a bespoke telemetry system integrated with version control, the Integrated Development Environments, and the Operating System. We validate DAT through observational studies, surveys, visualizations, and descriptive statistics. At Meta, DAT elevates the internal tool development workflow to the scientifically grounded, experiment-driven development flow already present for external-facing products. As such, DAT enables rigorous experimentation on long-standing software engineering questions like “do types make development more efficient?” DAT has powered experiments and case studies on more than 20 projects at Meta. Here, we highlight (1) a controlled experiment on introducing mock types, which showed that typed mocks in tests yield a 14% DAT improvement, (2) a case study on the development of automatic memoization in the React compiler, a 33% improvement, and (3) a framework to estimate the thousands of DAT hours saved annually through code sharing (>50>50% improvement). Overall, this paper contributes a novel, yet straightforward way to measure development velocity, DAT. It describes how the use of this metric has enabled rigorous experimentation on development productivity at industry scale, and paves the way toward a more scientific mindset in internal product development.
@@
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE-SEIP ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3786583.3786901isbn: 979-8-4007-2426-8/2026/04ccs: Software and its engineeringccs: Software and its engineering Software development methods
@@ 1. Introduction
For nearly two decades, Continuous Experimentation (CE) has been integral to external-facing product development in Big Tech (Kohavi et al., 2009; Tang et al., 2010; Karrer et al., 2021): Engineers frequently launch A/B experiments to evaluate the effects of new features. CE exposes a typically smaller cohort of users (A) to the proposed change, while withholding it from a control group (B), ensuring observed differences between A and B occur solely due to the proposed change and are free of interference. With a large user base, even small effect sizes can thus be detected. Only when a change is shown to have a statistically significant positive effect on the outcome measure, will it generally roll out (to B, too). In contrast to this rigorous, science-inspired development process for external products, is the development process of internal tools at the same companies often much less stringent: it mostly relies on a combination of basic usage numbers and qualitative feedback such as expert opinion, user surveys, anecdotes, and simply gut feel (Forsgren et al., 2024). This is puzzling given the tens of thousands of internal users and high value of engineering time, which constitute major operational expenses to the businesses. To improve internal development, a scientific, metrics- and experimentation-driven culture shift for internal development is needed. However, assessing the ultimate topline of internal development, measuring productivity accurately, remains challenging (Jaspan and Sadowski, 2019; Murphy-Hill et al., 2019; Beller et al., 2020), as traditional metrics are often inaccurate and easily gameable.
In this paper, we propose a possible pragmatic solution for such a productivity metric and how it can change an entire organization’s view on its development culture. We introduce Diff Authoring Time (DAT) as a key metric for assessing development velocity. DAT measures the active work time it takes engineers to develop a change, using a telemetry system integrated with version control, the Integrated Development Environment (IDE), ancillary tools, and the operating system. We validate DAT through observational studies, surveys, visualizations, and statistics. Out of more than 20 studies completed with DAT, we highlight three, demonstrating DAT’s impact across Meta:
@@
    A controlled experiment found that the introduction of mock types led to a 14% DAT improvement, marking a first in linking language-level features to productivity gains.
@@
    A post-hoc case study showed quantification of productivity gains from a new feature in the React Compiler, auto-memoization, resulting in 33% DAT reduction.
@@
    An estimation based counterfactual model showed estimated efficiency gains from code sharing as a >50>50% improvement across frameworks, saving thousands of engineering hours annually.
Overall, the contributions of this paper are three-fold:
  1. (1)
    We propose a novel, yet straight-forward way to measure development velocity using a combination of readily available telemetry sources including version control and the IDE.
  2. (2)
    We describe how the use of this metric has enabled rigorous experimentation on development productivity at industry scale, yielding new research findings such as that types do indeed make developers move faster.
  3. (3)
    We argue for a more scientific mindset to internal product development, going beyond qualitative evidence.
2. Related Work & Background
Figure 1. Example illustration of the most relevant DAT-covered and uncovered activities for two hypothetical diffs D123 and D987, and a comparison with other related metrics Code Gestation Time (CGT) and Time Spent by Diff (TSD).
@@ 2.1. Continuous Experimentation
“Continuous Experimentation” (CE) refers to the practice of regularly running experiments to validate one’s hypotheses. While this approach has become widespread in large, software-intense organizations and has defined much of the engineering work in Big Tech (Kohavi et al., 2009), it is not ubiquitous (Mäntylä et al., 2022). Challenges to adoption include a lack of resources and knowledge on CE, particularly in startups.
@@
Melegati et al. establish a “CE maturity model” with five stages of adoption and implementation (Melegati et al., 2020). When our work commenced, the engineering culture regarding CE in Meta’s infrastructure organization was somewhat similar to a startup: while Meta as a whole was naturally aware of CE (stage 1), it required a cultural mind shift to embrace it for Infra (stage 2). An organic need driven by lower observed output combined with leadership mandates initiated this process. To instill the new mindset and work processes throughout the organization, a particularly rigorous validation was necessary (stage 3), convincing key stakeholders. We derived results by focusing on a generally accepted measure (DAT), and integrating it into Meta’s already validated experimentation infrastructure (stage 4). Dedicated data scientists performed much of this work, leading to fast productionization of these results, including publishing the achieved wins in this paper (stage 5+).
2.2. Developer Productivity Metrics
Productivity metrics in Software Engineering have an over 50 year long history (Brooks, 1974), with dozens of different metrics proposed. Prominent product metrics include as lines of code (Jones and Nelson, 1976) or change count (Meli, 2001). Traditionally, one core problem with many of these metrics is their lack of meaning on “actual productivity” or failing to take into account the complexity of some tasks (a one-liner bug fix can be more complex than creating a green-field system and may have a higher impact). As a solution, function point analysis (Symons, 1988; Abran and Robillard, 2002) has been suggested, which delegates judgment of the complexity to a third-party, supported by guidelines. This is costly, may still be inaccurate, and has come out of fashion, perhaps due to the short iteration cycles in modern development that often obviate planning. Instead, a stronger emphasis on the developer experience (“DevEx”) has emerged in industry (Fagerholm and Münch, 2012; Noda et al., 2023). At the same time, the scientific community has started to focus on (self-)perceived productivity (Meyer et al., 2017; Cheng et al., 2022).
@@ 2.2. Developer Productivity Metrics
Recent works such as SPACE (Forsgren et al., 2021) and DX Core4 (Noda et al., 2024) provided an umbrella framework on how to think about software development productivity at large, but defer the exact metric choices to their users. The chosen metrics then often focus on easily measurable aspects of the development process, rather than the complexity of the authoring. For example, output metrics such as lines of code or change count emphasize the result of development but fail to capture the effort and complexity involved in the authoring lifecycle. Review time metrics, while good for measuring collaboration and delayed reviews, are only a sub-part of the authoring process. Another concern arises from the combination of many metrics into a construct that is now hard to reason about and act on.
Existing productivity metrics are thus either gameable, subjective, impossible to obtain accurately, a poor proxy of productivity, too complex, or some combination of the above (Kitchenham, 1997; Petersen, 2011; Symons, 1988; Huijgens and Solingen, 2014). One scholar even suggest abandoning productivity measurement altogether (Ko, 2019). The underlying challenge is that productivity measurement for knowledge workers, of which software engineers are a prime example (Ford et al., 2017), is a fundamentally unsolved problem (Ramírez and Nembhard, 2004).
@@
In this paper, we adopt a pragmatic approach, focusing on accurate time measurement for authoring. “Time spent coding” best aligns with developers’ own productivity perceptions (Beller et al., 2020), is objective, and can be measured at scale. To the best of our knowledge, no such metric for authoring has been formalized yet, since it would need to encompass all the different tools, workflows, and idiosyncrasies involved in authoring code. DAT does exactly this by measuring the active time taken to develop a diff while enabling insight into the components that impact it, such as IDE features, libraries, and programming languages. For instance, feature experimentation using DAT enables teams to assess the impact of specific improvements on developer velocity.
@@
We propose DAT as a conceptually simple, objective metric that is not merely theoretical, but of high practical value. At Meta, DAT supersedes Time Spent by Diff (TSD) and Code Gestation Time (CGT), illustrated in Figure 1. The figure shows engineering-related activities for two diffs (D123 and D987) across different tools, and in subsequent rows, which activities each metric covers. The illustrated sequential development workflow starts with coding on D123 and ends with landing it, with various activities (among them, coding on the unrelated D987) interspersed. DAT accurately covers both diffs’ core engineering, specifically in IDEs, while it cannot (yet) capture off-screen sketching time. TSD averaged coding time in a given period by the number of diffs published, but cannot differentiate between diffs D123 and D987 in the same period, as it lacks this level of precision. Thus, TSD was unable to run diff-level experiments. CGT, a composite metric, measured time from coding start to diff landing, mixing active engineering with technical time to release, double-counting overlapping engineering time (for D987), complicating interpretation and lacked validation.
2.3. Development Flow at Meta
2.3.1. A diff: Meta’s pull request equivalent
A diff, e.g., D123, represents a single code change including a natural text title, summary, and test plan (Beller et al., 2023), similar to a pull request (Gousios et al., 2014). Almost all code-based changes end up as (ideally) self-contained diffs that can stack on top of each other. A typical way to make diffs is by creating commits in an IDE such as VS Code (Marcey, 2019), using Sapling, Meta’s integrated, platform-independent version control client for multiple version control systems (Goode and Agarwal, 2014; Durham Goode, 2022). However, there are a number of other modalities for making diffs, including bespoke tools (Murali et al., 2024). Code review of all diffs happens on a custom version of the open-source Phabricator tool (Kudrjavets et al., 2022; Beller et al., 2023).
2.3.2. Diff-grain productivity measurements
Many productivity experiments make sense on the diff level, meaning that one developer can appear both in the A and B group for different diffs, naturally removing one of the biggest confounders. For example, a diff-level experiment could be to test the DAT effects of a new Continuous Integration (CI) signal, which, as one among many hundreds of CI signals, would not influence the developers’ core workflow or even be noticed if absent in the future. Similarly, switching out the base language model from LLaMA-33B to LLaMA-65B (Touvron et al., 2023) would likely also be somewhat opaque operation to a developer. Other experiments, though, need to be on the cohort level—it makes little sense and is practically impossible to switch one developer back-and forth between two VS Code versions with completely different user interfaces, for example. A diff-based measurement is still essential in these cases, as the diff level provides normalization for many other confounders such as length, complexity, language, framework, or testing. Overall, these characteristics make diffs a natural and high-coverage measurement unit for productivity and experimentation purposes.
@@ 3.1. Architecture
Diff Authoring Time (DAT) aims to capture all active human developer time spent on a particular diff, split by the different roles involved (e.g., author and reviewer). As Figure 1 shows, the key component of DAT is accurate representation of time spent in the IDE. DAT builds upon this to cover all diff-related activities and tools. Naturally, time off the computer, such as sketching, whiteboarding, or meeting time, will always be difficult to track.
@@
While there is an ongoing scientific debate on humans’ ability to multitask (Koch et al., 2018; Skaugset et al., 2016), active human work on a computer tends to be highly sequential, even if such individual sequences can be extremely short (e.g., rapid application switching often happens in <1s<1s): there can only be one application in focus at a given time, and such activity is atomically related to one diff only.
@@
This gives DAT advantageous properties that lend themselves to validation: it is non-overlapping (per user), meaning that two diffs by the same author must not have any overlapping time. As a corollary, DAT cannot exceed 24 hours per day (though it is, in practice, much shorter), can sensibly be aggregated (as opposed to CGT), and the sum of DAT in one day is bounded by TSD. Thus, DAT is a high-confidence, unobjectionable floor value of development time spent that is attributable to a diff. In fact, the comparison with TSD can reveal (1) improvement potential for DAT’s matching algorithm in cases where time is part of the larger TSD, but should also be part of DAT, or (2) potential for overall efficiency improvement, as coding-related activities in TSD correctly uncaptured in DAT might not have resulted in a productive output artifact (a diff), thus potentially representing “wasteful” effort. The goal for DAT is to be used purely as a productivity metric for tooling and workflows, not for evaluating people. Access and privacy restrictions ensure it can never be used in performance reviews, for example.
3.2. Algorithm
DAT consists of a main algorithm, precise matches, and a number of heuristics to capture additional time.
3.2.1. Core DAT algorithm: Precise matches
Figure 2. DAT’s core commit matching algorithm.
@@ 3.2.1. Core DAT algorithm: Precise matches
As the first two rows in Figure 1 depict, the core authoring of a diff starts with the initial trace of an activity that we can connect with this diff. Typically, this is the start of the first IDE session—most often in VS Code—in which the first commit is made that is then later connected to that diff. DAT includes only the time the engineer spends actively working on this diff, as the core DAT algorithm priorities precise, high-confidence matching—it omits IDE activity that does not result in a commit, or activity for commits that do not result in a diff. It only operates on tools that allow such high confidence attribution: IDEs and the Phabricator Diff review tool.
We obtain basic activity and foreground application (the application window in focus) through OS-level telemetry. We then correlate it to commit events using telemetry from the Sapling plugin inside the IDE. For authors of a diff, we only consider modifying commit events (e.g., commit or amend); for reviewers, we only consider non-modifying events (e.g., a checkout). Meta’s version control generates a new, unique commit hash even when modifying an existing commit and it does not allow rewriting of history. The Sapling plugin in the IDE talks to a host-wide Sapling server to which all Sapling client calls are routed. This has the advantage of tracking the underlying repository as a singleton, meaning that any interaction with it, be it from inside or outside the IDE (e.g., also from the command line), is captured.
@@
Figure 2 depicts the fundamental principle of DAT’s precise diff-to-session matching algorithm using the example of two diffs D123 and D987 that are developed sequentially in the same VS Code session: on top, the figure shows human activity as work sessions, along with point-in-time interactions with version control. It also shows the current head commit (i.e., the state of the repository from sl status, similar to git status). The figure depicts the principle that any commit $CH_{x}$ is attributed on the time axis to the IDE session $s_{x-1}$ that precedes it. Thus, at its core, the DAT algorithm performs a 1-left-shift of the underlying commits with regard to the telemetry sessions. For example, the commit $CH1$ made at time $t2$ is attributed to the VS Code session $s1$ that happened from $t1$ to $t2.$ This is because any change developed starting from $t1$ until $t2$ will become part of the commit $CH1$ made at time $t2$. Similarly, both $s2$ and $s4$ are mapped to $CH2$. $s2$ is interrupted by the offline activity $s3$ (“sketching”), which does not count toward DAT as OS-level telemetry indicates inactivity (e.g., no mouse movements or key input due to a coffee break). Timeout parameters have been chosen in line with existing work (Beller et al., 2017).
There are two corner cases:
  1. (1)
@@
    Any time a developer opens VS Code, an initial automatic checkout happens ($CH0$ in Figure 2). As the developer does not modify this pre-existing commit, the automatic commit checkout event filtered away by the DAT algorithm. Similarly, DAT removes all manual, non-modifying checkouts (e.g., the later checkout of the pre-existing commit $CH7$) via explicit filtering. The commit $CH8$, which becomes $CH7$’s successor, is attributed to D987.
  2. (2)
    For the “trailing session” $s7$, there is no commit mapping because any work done there would not actually be saved as a commit.
@@
We now have a precise commit-to-session mapping (hence the name “precise matches”), but not yet a mapping to diffs. This is trivial to achieve, as version control provides us with a list of all commits that make up a diff.
3.2.2. Additional DAT Heuristic: Anchor-DAT